### v1

# Towards Convexity in Anomaly Detection: A New Formulation of SSLM with Unique Optimal Solutions

Abstract
An unsolved issue in widely used methods such as Support Vector Data Description (SVDD) and Small Sphere and Large Margin SVM (SSLM) for anomaly detection is their nonconvexity, which hampers the analysis of optimal solutions in a manner similar to SVMs and limits their applicability in large-scale scenarios. In this paper, we introduce a novel convex SSLM formulation which has been demonstrated to revert to a convex quadratic programming problem for hyperparameter values of interest. Leveraging the convexity of our method, we derive numerous results that are unattainable with traditional nonconvex approaches. We conduct a thorough analysis of how hyperparameters influence the optimal solution, pointing out scenarios where optimal solutions can be trivially found and identifying instances of ill-posedness. Most notably, we establish connections between our method and traditional approaches, providing a clear determination of when the optimal solution is unique—a task unachievable with traditional nonconvex methods. We also derive the $\nu$-property to elucidate the interactions between hyperparameters and the fractions of support vectors and margin errors in both positive and negative classes.
Keywords: support vector data description, small sphere and large margin, anomaly detection, $\nu$-property, convex optimization
1 Introduction and Background
Anomaly detection, also known as outlier detection, is a crucial task in data analysis and aims to recognize whether a new observation belongs to the same distribution as existing observations. It is applicable across a wide range of domains where detecting anomalous data patterns is crucial, including anomaly detection, quality control, fraud detection, image analysis, and remote sensing. In these applications, only large datasets of the ‘normal’ or ‘positive’ condition (also known as positive examples, labeled as $y_{i}=+1$) are encountered, and the goal is to classify data into this primary class while not explicitly modeling or considering other classes of ‘abnormalities’ (also known as negative examples, labeled as $y_{i}=-1$).
A wide array of anomaly detection methods exists, and the support vector machine (SVM) is one of the most popular tools for this purpose. Several “(hyper)sphere-based” models have emerged among the various SVM-based approaches, each proving efficient and powerful in anomaly detection. These include the ball volume minimization model proposed by Schölkopf et al. (1995), the support vector data description (SVDD) Tax and Duin (2004), the soft-boundary SVDD Tax and Duin (1999b), as well as its convexified variant introduced by Chang et al. (2014), and the small sphere and large-margin SVM (SSLM) Wu and Ye (2009). At the core of these methods lies the construction of a hypersphere that encapsulates normal instances while ensuring that most, if not all, of the normal data points lie within or on the boundary of the hypersphere. This often leads to quadratically constrained quadratic programming (QCQP) problems involving the radius and center point of the hypersphere. These methods, particularly SVDD and SSLM, offer numerous advantages. They exhibit robustness when dealing with high-dimensional data, can handle nonlinear relationships through kernelization, and demonstrate a strong ability to generalize well to unseen data. Furthermore, they are valued for their interpretability, as anomalies are identified based on their distance from the center of the hypersphere.
Despite their wide application and numerous advantages, these methods also come with limitations, one of which stems from their nonconvex formulation, which is the central focus of this paper. Early models like Schölkopf et al. (1995) assume that normal data are encapsulated within a single hypersphere, and the optimal hypersphere is simply the one with the smallest volume, resulting in a convex model. The optimal hypersphere global can be easily determined in this case. However, real applications rarely accommodate this assumption, and such a model may become highly sensitive when dealing with outliers, non-convex anomalies, or data with complex structures. Modern models such as SVDD and SSLM allow negative examples and the misclassification of both positive and negative examples. Although the optimal solutions of such models possess powerful interpretation and generalization abilities, they give rise to nonconvex QCQP problems. This is a well-known critical issue that remains unsolved, with many associated issues and limitations stemming from it.
1.1 Related work
Many researchers have acknowledged the prevailing challenge posed by the nonconvexity issue inherent in sphere-based Support Vector Machines (SVMs). Over the past decades, several efforts have been devoted to addressing this intricate problem. Their focus has been centered on developing innovative strategies and algorithms to overcome the limitations imposed by nonconvexity, in a bid to enhance the performance and efficiency of sphere-based SVMs.
Consider the training data ${\bm{x}}_{1},\cdots,{\bm{x}}_{\ell}\in\mathcal{X},$ where for each $i\in\{1,\cdots,\ell\}$ (in compact notation: $i\in[1,\ell]$) the vector ${\bm{x}}_{i}$ represents the features, $\ell\in\mathbb{N}$ is the number of observations and $\mathcal{X}$ is the data domain and is assumed to be a compact subset of ${\mathbb{R}}^{N}$. Let $\Phi$ be a feature map $\mathcal{X}\to F$ such that the inner product in the image of $\Phi$ can be computed by evaluating a simple kernel $k({\bm{x}},{\bm{y}})=\langle\Phi({\bm{x}}),\Phi({\bm{y}})\rangle$ and denote $k_{ij}=k({\bm{x}}_{i},{\bm{x}}_{j})$. Denote ${\bm{a}}$ and $R$ for the ball center and radius, respectively. Schölkopf et al. (1995) minimized the volume $R^{2}$ of the ball while requiring all examples to be contained in this ball $\|{\bm{x}}_{i}-{\bm{a}}\|^{2}\leq R^{2},i\in[1,\ell]$. A test point is classified as normal if it is enclosed by this hypershere and abnormal otherwise.
However, this model is very sensitive to outlier data points. If there are one or several outlier data points in the training set, a very large hypersphere will be obtained. To accommodate the existence of outliers in the training data, Tax and Duin (1999a, b) proposed soft-boundary-SVDD:
$\begin{array}[]{cl}\mathop{{\mathrm{minimize}}}\limits_{{\bm{a}}\in{\mathbb{R}}^{N},R\in{\mathbb{R}},{\bm{\xi}}\in{\mathbb{R}}^{\ell}}&R^{2}+C\sum_{i=1}^{\ell}\xi_{i}\\ {\mathrm{subject\ to}}&\|{\bm{x}}_{i}-{\bm{a}}\|^{2}\leq R^{2}+\xi_{i},\xi_{i}\geq 0,i\in[1,\ell],\end{array}$ (1.1)
where $\xi_{i}$ is the relaxed variable and the positive parameter $C$ controls the trade-off between the volume and the empirical errors.
The nonconvexity term $R^{2}$ in the constraints of these methods can be removed by introducing a new nonnegative variable $r=R^{2}$. Chang et al. (2014) pointed out that the constraint $r\geq 0$ in this problem is not necessary for $C>1/\ell$, meaning in this case the optimal solution to the relaxation problem
$\begin{array}[]{ll}\mathop{{\mathrm{minimize}}}\limits_{{\bm{a}}\in F,r\in{\mathbb{R}},{\bm{\xi}}\in{\mathbb{R}}^{\ell}}&r+C\sum_{i=1}^{\ell}\xi_{i}\\ {\mathrm{subject\ to}}&\|{\bm{x}}_{i}-{\bm{a}}\|^{2}\leq r+\xi_{i},\xi_{i}\geq 0,i\in[1,\ell]\end{array}$ (1.2)
always has $r\geq 0$. If $C\leq 1/\ell$, then ${\bm{a}}_{*}=\frac{1}{\ell}\sum_{i=1}^{\ell}{\bm{x}}_{i}$ and $r_{*}=0$ is an optimal solution for (1.2). Notice that (1.2) is a convex problem and satisfies the Slater condition, ensuring strong duality. The dual problem of (1.2) is
$\begin{array}[]{ll}\mathop{\rm maximize}\limits_{{\bm{\alpha}}\in{\mathbb{R}}^{\ell}}&\sum_{i=1}^{\ell}\alpha_{i}{\bm{x}}_{i}^{T}{\bm{x}}_{i}-\sum_{i=1}^{\ell}\sum_{j=1}^{\ell}\alpha_{i}\alpha_{j}{\bm{x}}_{i}^{T}{\bm{x}}_{j}\\ {\mathrm{subject\ to}}&\sum_{i=1}^{\ell}\alpha_{i}=1,\\ &0\leq\alpha_{i}\leq C,i\in[1,\ell].\end{array}$ (1.3)
(Wang et al., 2011, Theorem 2) and (Chang et al., 2014, Theorem 2) showed that the optimal sphere center is always unique for the given dataset. As for the optimal ball radius, (Wang et al., 2011, Theorem 3) claimed that the sphere radius $r_{*}$ is not the unique optimal radius if and only if
$|{\mathcal{F}}_{r_{*}}|C=1$ (1.4)
where ${\mathcal{F}}_{r_{*}}=\{i:y_{i}=1,\|{\bm{x}}_{i}-{\bm{a}}_{*}\|^{2}-r_{*}>0\}$ is the examples that lie outside the sphere. Furthermore, (Chang et al., 2014, Theorem 5) showed that $r$ is optimal for (1.2) if and only if
$\max_{i:\alpha_{i}^{*}<C}\|{\bm{x}}_{i}-{\bm{a}}_{*}\|^{2}\leq r\leq\min_{i:\alpha_{i}^{*}>0}\|{\bm{x}}_{i}-{\bm{a}}_{*}\|^{2}.$ (1.5)
Suppose ${\bm{\alpha}}^{*}$ is an optimal solution for the dual problem (1.3). (Wang et al., 2011, Theorem 4) claimed that the radius of the ball is unique if and only if there exists $i_{0}\in\{1,\cdots,\ell\}$ such that $\alpha_{i_{0}}^{*}\in(0,C)$.
To separate the data set from the origin, Schölkopf et al. (1999, 2001) proposed one-class support vector machine (OCSVM):
$\begin{array}[]{cl}\mathop{{\mathrm{minimize}}}\limits_{{\bm{w}}\in F,{\bm{\xi}}\in{\mathbb{R}}^{\ell},\rho\in{\mathbb{R}}}&\frac{1}{2}\|{\bm{w}}\|^{2}-\rho+\frac{1}{\nu\ell}\sum_{i=1}^{\ell}\xi_{i}\\ {\mathrm{subject\ to}}&\langle{\bm{w}},\Phi({\bm{x}}_{i})\rangle\geq\rho-\xi_{i},\xi_{i}\geq 0,i\in[1,\ell],\end{array}$ (1.6)
where $\nu\in(0,1]$ is a given parameter. Note that although (1.2) is an equivalent convex formulation of (1.1), it remains the quadratic constraints. It is worthwhile to ask whether a convex quadratic program (QP) formulation for (1.1), as the well-known support vector machines (SVM)
$\begin{array}[]{cl}\mathop{{\mathrm{minimize}}}\limits_{{\bm{w}}\in F,{\bm{\xi}}\in{\mathbb{R}}^{\ell},\rho\in{\mathbb{R}}}&\frac{\nu}{2}\|{\bm{w}}\|^{2}+\frac{1}{\ell}\sum_{i=1}^{m}\xi_{i}+\frac{b}{\ell}\sum_{i=m+1}^{\ell}\xi_{i}\\ {\mathrm{subject\ to}}&y_{i}(\langle{\bm{w}},\Phi({\bm{x}}_{i})\rangle+b)\geq 1-\xi_{i},\xi_{i}\geq 0,i\in[1,\ell],\end{array}$ (1.7)
One can see (1.6) is a natural extension of (1.7) to the case of unlabeled data.
If negative examples (objects which should be rejected) are available, they can be incorporated into the training set to improve the description. For convenience, we assume that target objects are labeled $y_{i}=1$ for $i\in[1,m]$ and outlier objects are labeled $y_{i}=-1$ for $i\in[m+1,\ell]$. Tax and Duin (2001, 2004) extended the SVDD to learn a tight hyperball that encloses the majority of the normal data points while excluding the anomalies:
$\begin{array}[]{cl}\mathop{{\mathrm{minimize}}}\limits_{{\bm{a}}\in F,R\in{\mathbb{R}},{\bm{\xi}}\in{\mathbb{R}}^{\ell}}&R^{2}+C\sum_{i=1}^{m}\xi_{i}+D\sum_{i=m+1}^{\ell}\xi_{i}\\ {\mathrm{subject\ to}}&\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}\leq R^{2}+\xi_{i},\xi_{i}\geq 0,i\in[1,m]\\ &\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}\geq R^{2}-\xi_{i},\xi_{i}\geq 0,i\in[m+1,\ell],\end{array}$ (1.8)
where positive parameters $C$ and $D$ together control the trade-off between the volume and the empirical errors. These constraints enforce that the negative examples are located outside the hyperball, effectively pushing the boundary away from the anomalies. The addition of negative examples allows the model to better capture the boundary between normal and abnormal instances, potentially leading to improved performance in scenarios where anomalies are more diverse or prevalent. Wu and Ye (2009) proposed SSLM approach, which aims to define a hyperball that tightly encloses the majority of normal data points while maintaining a large margin from the decision boundary to separate normal data from outliers. The idea is to create a compact representation of positive examples while maximizing the separation between the positive and negative instances to achieve robust anomaly detection performance:
$\begin{array}[]{cl}\mathop{{\mathrm{minimize}}}\limits_{{\bm{a}}\in F,R,\rho\in{\mathbb{R}},{\bm{\xi}}\in{\mathbb{R}}^{\ell}}&R^{2}-\bar{\nu}\rho^{2}+\frac{1}{\bar{\nu}_{1}m}\sum_{i=1}^{m}\xi_{i}+\frac{1}{\bar{\nu}_{2}n}\sum_{i=m+1}^{\ell}\xi_{i}\\ {\mathrm{subject\ to}}&\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}\leq R^{2}+\xi_{i},\xi_{i}\geq 0,i\in[1,m]\\ &\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}\geq R^{2}+\rho^{2}-\xi_{i},\xi_{i}\geq 0,i\in[m+1,\ell],\end{array}$ (1.9)
where $n=\ell-m$, and $\rho^{2}\geq 0$ is the margin between the surface of the hyperball and the outlier training data, ${\bm{\xi}}\in{\mathbb{R}}^{\ell}$ is the vector of slack variables, and the positive parameter $\bar{\nu}$ controls the trade-off between the volume and margin, positive $\bar{\nu}_{1}$ and $\bar{\nu}_{2}$ control the trade-off between the volume and the errors. (1.9) can result in a closed and tight boundary around the normal data. Wu and Ye (2009) given the dual problem of (1.9):
$\begin{array}[]{ll}\mathop{\rm maximize}\limits_{{\bm{\alpha}}\in{\mathbb{R}}^{\ell}}&-\sum\limits_{i=1}^{\ell}\sum\limits_{j=1}^{\ell}y_{i}y_{j}k_{ij}\alpha_{i}\alpha_{j}+\sum\limits_{i=1}^{\ell}y_{i}k_{ii}\alpha_{i}\\ {\mathrm{subject\ to}}&\sum\limits_{i=1}^{\ell}y_{i}\alpha_{i}=1,\quad\sum\limits_{i=1}^{\ell}\alpha_{i}=2\bar{\nu}+1,\\ &0\leq\alpha_{i}\leq\tfrac{1}{\bar{\nu}_{1}m},i\in[1,m];\quad 0\leq\alpha_{i}\leq\frac{1}{\bar{\nu}_{2}n},i\in[m+1,\ell].\end{array}$ (1.10)
The term $R^{2}$ in the constraints of (1.8) and (1.9) associated with the positive examples can be removed by introducing new nonnegative variable $r=R^{2}$. Similarly, the term $\rho^{2}$ in the objective of (1.9) can be removed by introducing new nonnegative variable $t=\rho^{2}$. While the nonconvexity term $\|{\bm{a}}\|^{2}$ appearing in the constraint with negative examples in (1.8) and (1.9) is inherent, which results in a nonconvex QCQP problems.
1.2 Motivation
Nonconvexity is a prevalent issue in hypersphere-based SVM models, particularly in well-known methods such as SVDD and SSLM. This poses challenges in model property analysis, hyperparameter selection, and algorithm design.
Model Property Analysis. The nonconvexity of hypersphere-based SVMs significantly complicates the analysis of their solutions, especially when compared to hyperplane-based SVMs. Fundamental questions, such as whether a unique solution exists, remain unresolved. A key issue is the lack of strong duality, which is essential for working effectively in kernel space. Nonconvexity can result in an incorrect Lagrangian dual problem, maintaining a positive duality gap between the primal and dual solutions. Moreover, solving nonconvex problems to global optimality is often computationally difficult, if not impossible, and the solutions obtained by solvers are typically only local optima, limiting their usefulness for analysis.
Hyperparameter Selection. One of the most pressing issues with nonconvex models is the challenge of hyperparameter tuning. Parameters such as the sphere size, margin, and kernel function directly impact performance and generalization ability. The difficulty arises from the inability to guarantee “global” minima for each hyperparameter, as local minima make it hard to assess whether a chosen value is truly optimal. Additionally, balancing the size of the hypersphere (which represents normal data) against the margin separating outliers can be problematic, as local minima increase the sensitivity to outliers, resulting in suboptimal anomaly detection performance.
Algorithm Design. Algorithm design for hypersphere-based SVMs has been underexplored, particularly when compared to the more established hyperplane-based SVMs. For instance, Chapelle Chapelle (2007) demonstrated that the primal problem of SVMs can be efficiently solved, and Shalev-Shwartz et al. (2011) proposed the Pegasos algorithm, which uses stochastic subgradient descent to solve the primal problem of SVMs. We believe that resolving the nonconvexity issue in hypersphere-based SVMs would pave the way for the development of similarly powerful algorithms.
Approximation Methods and Efficiency. Numerous studies have proposed approximation methods to address the nonconvexity of SVDD and SSLM. For example, Chen et al. (2015) approximated SVDD using a Gaussian radial basis function kernel, facilitating easier solution acquisition. Wang and Xu (2023) replaced the hinge loss in SSLM with a ramp loss, converting the nonconvex problem into a difference-of-convex form, which was then solved using the concave-convex procedure. Guo et al. (2020) incorporated incremental learning into SSLM to improve training efficiency.
Nonconvexity also hampers many techniques designed to enhance computational efficiency during training. SSLM, in particular, involves multiple variables that cannot be explicitly represented as a linear combination of training samples. As a result, researchers Cao et al. (2020); Wang and Xu (2022); Yuan and Xu (2021) have focused on developing screening rules to identify and eliminate non-support vectors, reducing the size of the nonconvex problem. The first such rule, proposed by Cao et al. (2020), integrates the $\nu$-property, KKT conditions, and variational inequalities specific to SSLM.
This ongoing research demonstrates that addressing the nonconvexity in hypersphere-based SVMs can lead to significant advances in both theoretical analysis and practical algorithm development.
1.3 Contribution
The motivations discussed above prompted us to propose a convex optimization formulation to obtain the global optimal solution for the SSLM SVM with given hyperparameters. This allows us to analyze our method in a manner similar to SVMs, addressing a longstanding unresolved issue in hypersphere-based SVM models like SSLM and SVDD.
  • Convex Reformulation: We introduce a new kernelized SSLM model that is convex for any set of hyperparameters. Specifically, we prove that for prescribed values of these hyperparameters, our model reduces to either a convex quadratic programming (QP) or linear programming (LP) problem. This ensures that the global optimal solution can be found efficiently using existing large-scale QP/LP solvers.
  • Hyperparameter and Solution Interaction: We analyze the relationship between hyperparameters and the resulting optimal solutions, identifying “inappropriate” hyperparameter values that lead to ill-posed problems (where the optimal margin and objective value tend to infinity). Additionally, we highlight “trivial” hyperparameter choices for which the optimal solution can be explicitly determined. This analysis provides valuable guidance for hyperparameter selection.
  • Uniqueness and Dual Characterization: We demonstrate the uniqueness of the optimal solution for relevant hyperparameter values and characterize these solutions using dual variables. This type of analysis is rarely possible in nonconvex models, highlighting a key advantage of our convex approach.
  • $\nu$-Property: We derive the $\nu$-property of our model, describing the relationship between hyperparameters and the fractions of support vectors and margin errors in both the positive and negative classes. This property allows for flexible control of the SVM model’s complexity, which is particularly useful in situations such as class imbalance (common in anomaly detection) or when direct control over the number of support vectors is desired.
  • Connections to Traditional Methods: Our method establishes clear connections with traditional approaches like SVDD and SSLM, making it possible to apply the analyses and results presented here to those models. This is a significant advancement towards overcoming the limitations posed by nonconvexity in traditional hypersphere-based SVMs.
  • Numerical Performance: Through numerical experiments, we demonstrate that our convex method consistently outperforms traditional SSLM. While the nonconvexity of SSLM can lead to convergence to non-global minima when using local optimization algorithms, our convex approach reliably converges to the global minima, ensuring both efficiency and accuracy.
1.4 Outline
In Section 2, we introduce the convex SSLM model, and analyze the influence of hyperparameters on the optimal model. In Section 3, the connection of our proposed model to traditional SVDD and SSLM models is established. Section 4 presents the derivation of the Lagrangian dual problem for the proposed model, proves the uniqueness of optimal solutions for relevant hyperparameter values, and characterizes the optimal solution using dual variables, highlighting the $\nu$-property of our convex QP formulation. In Section 5, we provide experimental results that demonstrate the advantages of our method over traditional SSLM. Finally, Section 6 concludes the paper.
2 The Convex SSLM
In this section, we present our convex SSLM (CSSLM) on the data set $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{\ell}$ with $y_{i}=+1,i\in[1,m]$ and $y_{i}=-1,i\in[m+1,\ell]$ with $\ell:=m+n$. The $m$ positive examples are required to be contained in the small hypersphere with radius $r\geq 0$,
$\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}\leq r,\quad y_{i}=+1.$
Let $t\geq 0$ be the margin between the two spheres. The $n$ negative examples are excluded from the hypersphere
$\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}\geq r+t,\quad y_{i}=-1.$
The empirical risk is given by
$\varphi_{i}({\bm{a}},r,t)=\left\{\begin{array}[]{rl}\Big{(}\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}-r\Big{)}_{+},&y_{i}=1,\\ b\Big{(}r+t-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}\Big{)}_{+},&y_{i}=-1,\end{array}\right.$
with $(z)_{+}:=\max\{z,0\}$. By minimizing the small sphere and the empirical risk, we arrive at the regularized empirical risk minimization formulation of (1.9)
$\inf\limits_{{\bm{a}}\in F,r\geq 0,t\geq 0}\ g({\bm{a}},r,t):=\tfrac{1}{2}(\nu r-\mu t)+\tfrac{1}{2\ell}\sum\limits_{i=1}^{\ell}\varphi_{i}({\bm{a}},r,t).$ ($\mathcal{F}$)
Specifically, $g$ takes the form
$g({\bm{a}},r,t)=\tfrac{1}{2}(\nu r-\mu t)+\tfrac{1}{2\ell}\sum\limits_{i=1}^{m}(\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}-r)_{+}+\tfrac{b}{2\ell}\sum\limits_{i=m+1}^{\ell}(r+t-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2})_{+}.$
Here prescribed $\nu>$ controls the trade-off between the volume and the loss, $\mu>0$ controls the trade-off between the volume and the margin, and $b\geq 1$ controls the trade-off between the loss by positive examples and that by negative examples.
Before proceeding to propose the CSSLM based on ($\mathcal{F}$), we first point out in the following (See Appendix A for the proof) that the objective of ($\mathcal{F}$) is unbounded below if the value of $\mu$ is set too large.
Lemma 1
If $\mu>\min\{m/{\ell},bn/{\ell}\}$, then ($\mathcal{F}$) is unbounded below.
Obviously, ($\mathcal{F}$) is a nonconvex model. We do not suggest directly solving an nonconvex model like ($\mathcal{F}$). In the rest of this paper, we assume $\nu$ and $\mu$ are selected such that
$\boxed{\mu\leq\min\{\tfrac{m}{\ell},\tfrac{bn}{\ell}\}.}$ (2.1)
We show that for all values satisfying (2.1), the global optimal solution of ($\mathcal{F}$) can be found by solving a convex problem.
First, suppose the parameters are selected such that $\nu+\mu<m/{\ell}$, the global optimal solution of ($\mathcal{F}$) can be found by solving the convex problem
$\inf\limits_{{\bm{a}}\in F,t\geq 0}\ g({\bm{a}},r,t).$ (2.2)
Compared with ($\mathcal{F}$), the absence of constraint $r\geq 0$ represents a significant difference, which enables (2.2) to be convex.
Theorem 2
If $\nu+\mu<m/{\ell}$, then the global optimal solution $({\bm{a}}_{*},r_{*},t_{*})$ of (2.2) satisfies $r_{*}\geq 0$ and therefore is global optimal for ($\mathcal{F}$). Conversely, every global minimizer $(\bar{\bm{a}},\bar{r},\bar{t})$ of ($\mathcal{F}$) is also optimal for (2.2).
Proof  Suppose $({\bm{a}}_{*},r_{*},t_{*})$ is the optimal solution of (2.2). Assume by contradiction that $r_{*}<0$. To derive a contradiction, we now investigate the value of $g$ in two cases based on the value of $r_{*}+t_{*}$. For this purpose, define
$\hat{{\bm{a}}}:=\arg\min_{{\bm{a}}\in F}\left\{\frac{1}{m}\sum_{i=1}^{m}\|\Phi(x_{i})-{\bm{a}}\|^{2}\right\}=\frac{1}{m}\sum_{i=1}^{m}\Phi(x_{i}).$ (2.3)
If $r_{*}+t_{*}\leq 0$, we have
$\begin{array}[]{ll}g({\bm{a}}_{*},r_{*},t_{*})&=\frac{1}{2}(\nu+\mu-\frac{m}{\ell})r_{*}-\frac{\mu}{2}(r_{*}+t_{*})+\frac{1}{2\ell}\sum_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\\ &\geq\frac{1}{2}(\nu+\mu-\frac{m}{\ell})r_{*}+\frac{1}{2\ell}\sum_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\\ &>\frac{1}{2\ell}\sum_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\\ &\geq\frac{1}{2\ell}\sum_{i=1}^{m}\|\Phi({\bm{x}}_{i})-\hat{\bm{a}}\|^{2}\\ &=g(\hat{{\bm{a}}},0,0),\end{array}$
where the first inequality follows from $r_{*}+t_{*}\leq 0$ and $\mu\geq 0$, and the strict inequality is by $\nu+\mu<m/{\ell}$ and $r_{*}<0$, and the last inequality is due to the fact that $\hat{\bm{a}}$ is the minimizer of $\sum_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}$ because (2.3). Obviously, $(\hat{\bm{a}},0,0)$ is feasible for (2.2). This contradicts the optimality of $({\bm{a}}_{*},r_{*},t_{*})$.
If $r_{*}+t_{*}>0$, we have
$g({\bm{a}}_{*},r_{*},t_{*})$
$=$ $\tfrac{1}{2}(\nu+\mu-\frac{m}{\ell})r_{*}-\tfrac{\mu}{2}(r_{*}+t_{*})+\tfrac{1}{2\ell}\sum\limits_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}+\tfrac{b}{2\ell}\sum\limits_{i=m+1}^{\ell}(r_{*}+t_{*}-\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2})_{+}$
$>$ $-\tfrac{\mu}{2}(r_{*}+t_{*})+\tfrac{1}{2\ell}\sum\limits_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}+\tfrac{b}{2\ell}\sum\limits_{i=m+1}^{\ell}(r_{*}+t_{*}-\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2})_{+}$
$=$ $g({\bm{a}}_{*},0,r_{*}+t_{*}),$
where the inequality follows by $\nu+\mu<m/{\ell}$ and $r_{*}<0$. Notice that $({\bm{a}}_{*},0,r_{*}+t_{*})$ is also feasible for (2.2) by $r_{*}+t_{*}>0$. This contradicts the optimality of $({\bm{a}}_{*},r_{*},t_{*})$. Overall, we have shown that $r_{*}\geq 0$, so that $({\bm{a}}_{*},r_{*},t_{*})$ is also feasible for ($\mathcal{F}$). Hence, it is globally optimal for ($\mathcal{F}$).
Conversely, let $(\bar{\bm{a}},\bar{r},\bar{t})$ be globally optimal for ($\mathcal{F}$). It follows that $f(\bar{\bm{a}},\bar{r},\bar{t})\leq f({\bm{a}}_{*},r_{*},t_{*})$ because $({\bm{a}}_{*},r_{*},t_{*})$ is feasible for ($\mathcal{F}$); moreover, $(\bar{\bm{a}},\bar{r},\bar{t})$ is also feasible for (2.2). Hence, $(\bar{\bm{a}},\bar{r},\bar{t})$ is also optimal for (2.2).
As for solving (2.2), letting $k_{ij}=\langle\Phi({\bm{x}}_{i}),\Phi({\bm{x}}_{j})\rangle$, we introduce the intermediate variable
$s=\|{\bm{a}}\|^{2}-r,$ (2.4)
instead of using $r$ as the variable, so that the empirical risk becomes
$\psi_{i}({\bm{a}},s,t)=\left\{\begin{array}[]{rl}\Big{(}k_{ii}-2\langle\Phi({\bm{x}}_{i}),{\bm{a}}\rangle+s\Big{)}_{+},&y_{i}=1,\\ b\Big{(}-s+t-k_{ii}+2\langle\Phi({\bm{x}}_{i}),{\bm{a}}\rangle\Big{)}_{+},&y_{i}=-1,\end{array}\right.$
Now (2.2) reverts to
$\inf\limits_{{\bm{a}}\in F,t\geq 0}\ f({\bm{a}},s,t):=\tfrac{1}{2}\nu(\|{\bm{a}}\|_{2}^{2}-s)-\tfrac{1}{2}\mu t+\tfrac{1}{2\ell}\sum\limits_{i=1}^{m}\psi_{i}({\bm{a}},s,t).$ (2.5)
The radius $r$ of the sphere appears as a latent variable in this model, and the small sphere is attained by minimizing $\|{\bm{a}}\|^{2}-s$ instead of $r$. By introducing the auxiliary variable $\xi_{i}$, (2.2) is indeed a convex QP model [^1] Note that the objective function of (2.5) is a scaled one of ($\mathcal{P}$) with a factor $1/\ell$.:
$\begin{array}[]{cl}\mathop{\rm{\mathrm{minimize}}}\limits_{{\bm{a}}\in F,s,t\in{\mathbb{R}},{\bm{\xi}}\in{\mathbb{R}}^{\ell}}&h({\bm{a}},s,t,{\bm{\xi}}):=\frac{\ell\nu}{2}(\|{\bm{a}}\|^{2}-s)-\frac{\ell\mu}{2}t+\sum\limits_{i=1}^{m}\xi_{i}+b\sum\limits_{i=m+1}^{\ell}\xi_{i}\\ {\mathrm{subject\ to}}&y_{i}[\frac{k_{ii}}{2}-\langle\Phi({\bm{x}}_{i}),{\bm{a}}\rangle+\frac{s}{2}]\hskip 16.00008pt-\xi_{i}\leq 0,\ \xi_{i}\geq 0,\quad i\in[1,m]\\ &y_{i}[\frac{k_{ii}}{2}-\langle\Phi({\bm{x}}_{i}),{\bm{a}}\rangle+\frac{s}{2}]+\frac{t}{2}-\xi_{i}\leq 0,\ \xi_{i}\geq 0,\quad i\in[m+1,\ell]\\ &\hskip 133.00023ptt\geq 0.\end{array}$ ($\mathcal{P}$)
It is often more convenient for SVM methods to train in the kernel space without knowing the feature mapping $\phi$ explicitly, then transfer ($\mathcal{P}$) from the kernel space to the feature mapping $\Phi$. Then we have the convex QP problem
$\min\limits_{{\bm{\beta}}\in{\mathbb{R}}^{\ell},s\in{\mathbb{R}},t\geq 0}\frac{\nu}{2}({\bm{\beta}}^{T}{\bm{K}}{\bm{\beta}}-s)-\frac{\mu}{2}t+\frac{1}{\ell}\sum\limits_{i=1}^{m}\phi_{i}({\bm{\beta}},s,t)+\frac{b}{\ell}\sum\limits_{i=m+1}^{\ell}\phi_{i}({\bm{\beta}},s,t),$ ($\mathcal{K}$)
where
$\phi_{i}({\bm{\beta}},s,t)=\left\{\begin{array}[]{rl}\Big{(}\frac{k_{ii}+s}{2}-\sum_{j=1}^{\ell}\beta_{j}k_{ij}\Big{)}_{+},&y_{i}=1,\\ b\Big{(}\frac{t}{2}-(\frac{k_{ii}+s}{2}-\sum_{j=1}^{\ell}\beta_{j}k_{ij})\Big{)}_{+},&y_{i}=-1.\end{array}\right.$
Problem ($\mathcal{K}$) is suitable to incorporate with different kernel techniques. The sphere center in ($\mathcal{K}$) now also becomes a hidden variable. Nevertheless, it is trivial to recover the optimal solution to ($\mathcal{K}$) by the following representer theorem.
Theorem 3
Let $F$ be a reproduced kernel Hilbert space with positive definite symmetric kernel $k(\cdot,\cdot)$. Suppose $\nu+\mu<m/{\ell}$. Then $({\bm{a}}_{*},r_{*},t_{*})$ is optimal for (2.2) if and only if $({\bm{\beta}}_{*},s_{*},t_{*})$ is optimal for ($\mathcal{K}$) with $r_{*}=\|{\bm{a}}_{*}\|^{2}-s_{*}$ and
${\bm{a}}_{*}=\sum_{i=1}^{\ell}\beta_{i}^{*}\Phi({\bm{x}}_{i}).$ (2.6)
Proof  Define
${\bar{F}}_{\ell}:={\rm span}\{\Phi({\bm{x}}_{1}),\cdots,\Phi({\bm{x}}_{\ell})\}=\left\{{\bm{a}}\in F:{\bm{a}}(\cdot)=\sum_{i=1}^{\ell}\beta_{i}k({\bm{x}}_{i},\cdot),{\bm{\beta}}\in{\mathbb{R}}^{\ell}\right\}.$
We first show that if $({\bm{a}}_{*},r_{*},t_{*})$ is optimal for (2.2), then ${\bm{a}}_{*}\in{\bar{F}}_{\ell}$. By contradiction, suppose that ${\bm{a}}_{*}\notin{\bar{F}}_{\ell}$. Since ${\bar{F}}_{\ell}$ is a linear subspace of $F$, we can decompose ${\bm{a}}_{*}\in F$ as ${\bm{a}}_{*}=\bar{\bm{a}}_{*}+{\bm{a}}_{*}^{\perp}$ with $\bar{\bm{a}}_{*}\in{\bar{F}}_{\ell}$ and $\bm{0}\neq{\bm{a}}_{*}^{\perp}\in{\bar{F}}_{\ell}^{\perp}$. Moreover, since $\Phi({\bm{x}}_{i})\in\bar{F}_{\ell}$, the Pythagorean theorem then gives
$\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}=\|\Phi({\bm{x}}_{i})-\bar{\bm{a}}_{*}\|^{2}+\|{\bm{a}}_{*}^{\perp}\|^{2}\ {\rm for}\ i\in[1,\ell].$
Elementary computations show that
$g({\bm{a}}_{*},r_{*},t_{*})=g(\bar{\bm{a}}_{*},r_{*}-\|{\bm{a}}_{*}^{\perp}\|^{2},t_{*})+\frac{\nu}{2}\|{\bm{a}}_{*}^{\perp}\|^{2}.$
Thus
$g({\bm{a}}_{*},r_{*},t_{*})>g(\bar{\bm{a}}_{*},r_{*}-\|{\bm{a}}_{*}^{\perp}\|^{2},t_{*})$
because ${\bm{a}}_{*}^{\perp}\neq\bm{0}$. This violates the global optimality of $({\bm{a}}_{*},r_{*},t_{*})$ for (2.2). Hence ${\bm{a}}_{*}\in\bar{F}_{\ell}$.
Conversely, if $({\bm{\beta}}_{*},s_{*},t_{*})$ is optimal for ($\mathcal{K}$), $({\bm{a}}^{*},r_{*},t_{*})=(\sum_{i=1}^{\ell}\beta_{i}^{*}\Phi({\bm{x}}_{i}),\|{\bm{a}}_{*}\|^{2}-s_{*},t^{*})$ must be optimal for (2.2). Otherwise, there exists $(\hat{\bm{a}},\hat{r},\hat{t})\neq({\bm{a}}^{*},r_{*},t_{*})$ such that $g(\hat{\bm{a}},\hat{r},\hat{t})<g({\bm{a}}^{*},r_{*},t_{*})$. However, we have shown there exists a decomposition $\hat{\bm{a}}=\sum_{i=1}^{\ell}\hat{\beta}_{i}\Phi({\bm{x}}_{i})$ such that $(\hat{\bm{\beta}},\|{\bm{a}}\|_{2}^{2}-\hat{r},\hat{t})$ yields a smaller objective ($\mathcal{K}$)—a contradiction. Therefore, $({\bm{a}}_{*},r_{*},t_{*})$ is optimal for (2.2).
As we show in the next, for other choices of parameter values $\nu,\mu$ and $b$, although the constraint $r\geq 0$ is needed, the global optimal solution of (2.2) can still be found by solving convex problems. This result is summarized in the following theorem. The proof is provided in Appendix A.
Theorem 4
Suppose $\nu+\mu\geq m/{\ell}$ is satisfied.
$(i)$ If $\mu=0$, the infimum of ($\mathcal{F}$) is attained at
$\Omega:=\Big{\{}(\hat{\bm{a}},r,t):\Big{(}\nu-\frac{m}{\ell}\Big{)}r=0,0\leq r\leq\min_{y_{i}=1}\{\|\Phi({\bm{x}}_{i})-\hat{\bm{a}}\|^{2}\},r+t\leq\min_{y_{i}=-1}\{\|\Phi({\bm{x}}_{i})-\hat{\bm{a}}\|^{2}\},t\geq 0\Big{\}}.$
In particular, $r^{*}=0,t^{*}=\min\limits_{y_{i}=-1}\|\Phi({\bm{x}}_{i})-\hat{\bm{a}}\|^{2},{\bm{a}}^{*}=\hat{\bm{a}}$ is an optimal solution of ($\mathcal{F}$).
$(ii)$ If $0<\mu<m/{\ell}$, $({\bm{a}}^{*},z^{*})$ is optimal for
$\begin{array}[]{cl}\mathop{{\mathrm{minimize}}}\limits_{{\bm{a}}\in F,z\in{\mathbb{R}},{\bm{\xi}}\in{\mathbb{R}}^{n}}&\frac{\ell\lambda}{2}\|{\bm{a}}\|^{2}+\frac{\ell\mu}{2}z-m\langle\hat{\bm{a}},{\bm{a}}\rangle+b\sum\limits_{i=m+1}^{\ell}\xi_{i}\\ {\mathrm{subject\ to}}&-\frac{k_{ii}}{2}+\langle\Phi({\bm{x}}_{i}),{\bm{a}}\rangle-\frac{z}{2}-\xi_{i}\leq 0,\ \xi_{i}\geq 0,i\in[m+1,\ell]\end{array}$ (2.7)
with $\lambda=m/\ell-\mu$, then the solution satisfies $\|{\bm{a}}^{*}\|^{2}\geq z^{*}$, and $({\bm{a}}^{*},0,\|{\bm{a}}^{*}\|^{2}-z^{*})$ is optimal for ($\mathcal{F}$). Conversely, if $({\bm{a}}^{*},r^{*},t^{*})$ is optimal for ($\mathcal{F}$), $({\bm{a}}^{*},\|{\bm{a}}^{*}\|^{2}-(r^{*}+t^{*}))$ is optimal for (2.7).
$(iii)$ If $\mu=m/\ell$, $({\bm{a}}^{*},z^{*})$ is optimal for
$\begin{array}[]{cl}\mathop{{\mathrm{minimize}}}\limits_{a\in F,z\in\mathbb{R},\xi\in\mathbb{R}^{l}}&\frac{\ell\mu}{2}z-m\langle\hat{\bm{a}},{\bm{a}}\rangle+b\sum\limits_{i=m+1}^{l}\xi_{i}\\ {\mathrm{subject\ to}}&-\frac{k_{ii}}{2}+\langle\Phi({\bm{x}}_{i}),{\bm{a}}\rangle-\frac{z}{2}-\xi_{i}\leq 0,\ \xi_{i}\geq 0,i\in[m+1,\ell],\end{array}$ (2.8)
then the solution satisfies $\|{\bm{a}}^{*}\|^{2}\geq z^{*}$, and $({\bm{a}}^{*},0,\|{\bm{a}}^{*}\|^{2}-z^{*})$ is optimal for ($\mathcal{F}$). Conversely, if $({\bm{a}}^{*},r^{*},t^{*})$ is optimal for ($\mathcal{F}$), $({\bm{a}}^{*},\|{\bm{a}}^{*}\|^{2}-(r^{*}+t^{*}))$ is optimal for (2.8).
The entire scheme of the proposed CSSLM is described in the flowchart in Figure 1. The procedure starts with choosing the appropriate parameter values of $\mu,\nu$ and then decides which convex problems to solve, or the solution is trivially obtained. Figure 2 depicts the optimal value of $r^{*}$ dependent on different choices of $\mu$ and $\nu$.
Input $(\nu,\mu,b)$$\mu\leq\min\{m/{\ell},bn/{\ell}\}$return $\infty$- ∞$\mu+\nu<m/{\ell}$$0<\mu<m/{\ell}$$\mu=m/{\ell}$$\mu=0$solve QP ($\mathcal{P}$)solve QP (2.7)solve LP (2.8)return $(\hat{\bm{a}},0,0)$NoYesYesNo
Figure 1: The flowchart of CSSLM.
Figure 2: Optimal $r_{*}$ dependent of $\mu$ and $\nu$.
3 Connection to SSLM and SVDD
We have proved that our proposed method can find the global minimizer $({\bm{a}}_{*},r_{*},t_{*})$ of ($\mathcal{F}$) for different scenarios of parameters $\nu$, $\mu$ and $b$. Now, we show our proposed model can easily render the global optimal solutions of SSLM (1.9) and SVDD (including the one-class soft-margin SVDD (1.1) and the two-class soft-margin SVDD (1.8)).
Letting $r=R^{2}$ (or, $R^{2}=\|{\bm{a}}\|^{2}-s$ in $f({\bm{a}},s,t)$) and $t=\rho^{2}$, ($\mathcal{F}$) reverts to
$\min\limits_{{\bm{a}}\in F,R,\rho}\ \frac{\nu}{2}\left(R^{2}-\frac{\mu}{\nu}\rho^{2}+\frac{1}{\nu\ell}\sum\limits_{i=1}^{m}(d_{i}^{2}-R^{2})_{+}+\frac{b}{\nu\ell}\sum\limits_{i=m+1}^{\ell}(R^{2}+\rho^{2}-d_{i}^{2})_{+}\right)$ (3.1)
with $d_{i}=\|\Phi({\bm{x}}_{i})-{\bm{a}}\|$. It is the unconstrained version of the SSLM problem (1.9) with parameters chosen as
$(\bar{\nu},\bar{\nu}_{1},\bar{\nu}_{2})=\left(\frac{\mu}{\nu},\frac{\ell\nu}{m},\frac{\ell\nu}{nb}\right)$ (3.2)
where the objective of (3.1) is that of (1.9) scaled by $\nu/2$. Since (3.1) and (1.9) share the same global minimizers, then $({\bm{a}}_{*},R_{*},\rho_{*})$ is the global minimizer $({\bm{a}}_{*},R_{*},\rho_{*})$ of (1.9) with
$R_{*}=\sqrt{r_{*}}\text{ and }\rho_{*}=\sqrt{t_{*}}.$ (3.3)
On the other hand, setting $\mu=0,C=\frac{1}{\nu\ell}$ and $D=\frac{b}{\nu\ell}$ in (3.1) yields
$\begin{array}[]{cl}\min\limits_{{\bm{a}}\in F,R,\rho}\hat{g}({\bm{a}},R,\rho):=\frac{\nu}{2}\left(R^{2}+C\sum\limits_{i=1}^{m}(d_{i}^{2}-R^{2})_{+}+D\sum\limits_{i=m+1}^{\ell}[R^{2}+\rho^{2}-d_{i}^{2}]_{+}\right).\end{array}$ (3.4)
If $m=\ell$, then the last summation in the objective function of (3.4) disappears and (3.4) is obviously equivalent to the one-class soft-margin SVDD problem (1.1). Hence, $({\bm{a}}_{*},\sqrt{r_{*}})$ is the global minimizer of (1.1).
If $m<\ell$, we show in the following theorem that (3.4) is equivalent to the unconstrained version of the two-class soft-margin SVDD (1.8), which can be rewritten as
$\min\limits_{{\bm{a}}\in F,R}\ \hat{g}({\bm{a}},R,0).$ (3.5)
Theorem 5
Suppose $\mu=0$ and $m<\ell$. For any $({\bm{a}}_{*},R_{*},\rho_{*})$ optimal for (3.4), $({\bm{a}}_{*},R_{*})$ is globally optimal for (3.5). Conversely, for any $({\bm{a}}_{*},R_{*})$ optimal for (3.5), $({\bm{a}}_{*},R_{*},0)$ is globally optimal for (3.4).
Proof  For any optimal solution $({\bm{a}}_{*},R_{*},\rho_{*})$ for (3.4), $\hat{g}({\bm{a}}_{*},R_{*},\rho_{*})\geq\hat{g}({\bm{a}}_{*},R_{*},0)$ since $\hat{g}$ is increasing in $|\rho|$. Therefore, $({\bm{a}}_{*},R_{*},0)$ is also optimal for (3.4). It follows that $\hat{g}({\bm{a}},R,0)\geq\hat{g}({\bm{a}}_{*},R_{*},0)$ for any $({\bm{a}},R)$, meaning $({\bm{a}}_{*},R_{*})$ is global optimal for (3.5).
Conversely, if $({\bm{a}}_{*},R_{*})$ is optimal for (3.5), then
$\hat{g}({\bm{a}}_{*},R_{*},0)\leq\hat{g}({\bm{a}},R,0)\leq\hat{g}({\bm{a}},R,\rho)$
for any $({\bm{a}},R,\rho)$. Therefore, $({\bm{a}}_{*},R_{*},0)$ is optimal for (3.4).
Theorem 2 and Theorem 5 imply that when setting $\mu=0$, $m<\ell$ and $\nu<m/{\ell}$, our proposed CSSLM renders the global minimizer $({\bm{a}}_{*},\sqrt{r_{*}})$ of (1.8) by obtaining an optimal $({\bm{a}}_{*},r_{*},t_{*})$ to (2.2).
Remark 6
Theorem 5 does not imply that any optimal $\rho_{*}$ for (3.4) must be zero. In fact, if there exists $i_{*}\in[m+1,\ell]$ such that the optimal $({\bm{a}}_{*},R_{*},\rho_{*})$ for (3.4) satisfies
$\|\Phi({\bm{x}}_{i_{*}})-{\bm{a}}_{*}\|\leq R_{*},$ (3.6)
then $\hat{g}({\bm{a}}_{*},R_{*},0)<\hat{g}({\bm{a}}_{*},R_{*},\rho)$ for any $\rho>0$; therefore, the optimal solution satisfies $\rho_{*}=0$. On the other hand, if (3.6) is not satisfied, meaning
$\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|>R_{*},\quad\text{ for any }i\in[m+1,\ell],$
then $({\bm{a}}_{*},R_{*},\rho_{*})$ is globally optimal for (3.4) for any
$\rho_{*}\in[0,\hat{\rho}],\quad\text{with }\ \hat{\rho}:=\min_{i\in[m+1,\ell]}\{\|\Phi({\bm{x}}_{i_{*}})-{\bm{a}}_{*}\|-R_{*}\}.$
We summarize the connections between our proposed CSSLM and the state-of-the-art SSLM (1.9), one-class SVDD (1.1) and two-class SVDD (1.8) in Figure 3. Therefore, the global minimizers of SSLM and SVDD can be found by applying polynomial algorithms to solve CSSLM.
$({\bm{a}}_{*},r_{*},t_{*})$$\mu>0$$\mu=0$$({\bm{a}}_{*},\sqrt{r_{*}},\sqrt{t_{*}})$$({\bm{a}}_{*},\sqrt{r_{*}})$$m=\ell$$m<\ell$SSLMOne-class SVDDTwo-class SVDDGlobal optimal Global optimal Global optimal
Figure 3: Connection to SSLM and SVDD
4 Existence and Uniqueness of Optimal Solutions
In this section, we state and prove our main results on the existence and uniqueness of optimal solutions of the problems ($\mathcal{P}$), (2.7) and (2.8) involved in our proposed model.
4.1 Existence of optimal solutions
One possible and rather common approach for solving the convex optimization problem is to solve its dual problem. Elementary computation show that the Lagrangian dual of ($\mathcal{P}$) admits an explicit form, i.e.
$\begin{array}[]{ll}\mathop{\rm maximize}\limits_{{\bm{\alpha}}\in{\mathbb{R}}^{\ell}}&-\frac{1}{2\ell\nu}\sum\limits_{i=1}^{\ell}\sum\limits_{j=1}^{\ell}y_{i}y_{j}k_{ij}\alpha_{i}\alpha_{j}+\frac{1}{2}\sum\limits_{i=1}^{\ell}y_{i}k_{ii}\alpha_{i}\\ {\mathrm{subject\ to}}&\sum\limits_{i=1}^{\ell}y_{i}\alpha_{i}=\ell\nu,\quad\sum\limits_{i=m+1}^{\ell}\alpha_{i}\geq\ell\mu,\\ &0\leq\alpha_{i}\leq 1,i\in[1,m];\quad 0\leq\alpha_{i}\leq b,i\in[m+1,\ell].\end{array}$ ($\mathcal{D}$)
Theorem 7
Suppose $\mu+\nu<m/\ell$ holds. Problem ($\mathcal{P}$) and its dual problem ($\mathcal{D}$) are solvable and the strong duality holds at the primal-dual optimal solution. In particular, $({\bm{a}},s,t,{\bm{\xi}},{\bm{\alpha}})$ is the primal-dual optimal solution if and only if it satisfies
$\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}-r-2\xi_{i}\leq 0,\xi_{i}\geq 0,$ $\ \ i\in[1,m],$ (4.1a)
$q-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}-2\xi_{i}\leq 0,\xi_{i}\geq 0,$ $\ \ i\in[m+1,\ell],$ (4.1b)
${\bm{a}}=\frac{1}{\ell\nu}\sum_{i=1}^{\ell}y_{i}\Phi({\bm{x}}_{i})\alpha_{i},$ (4.1c)
$\sum_{i=1}^{\ell}y_{i}\alpha_{i}=\ell\nu,\,\,\sum_{i=m+1}^{\ell}\alpha_{i}\geq\ell\mu,$ (4.1d)
[…]
$\xi_{i}(1-\alpha_{i})=0,\ \ i\in[1,m];\quad\xi_{i}(b-\alpha_{i})=0,$ $\ i\in[m+1,\ell],$ (4.1f)
$(\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}-r-2\xi_{i})\alpha_{i}=0,$ $\ \ i\in[1,m],$ (4.1g)
$(q-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}-2\xi_{i})\alpha_{i}=0,$ $\ \ i\in[m+1,\ell],$ (4.1h)
$t\geq 0,t\big{(}\sum_{i=m+1}^{\ell}\alpha_{i}-\ell\mu\big{)}=0$ (4.1i)
with $r=\|{\bm{a}}\|^{2}-s$ and $q=r+t$.
Proof  Problem ($\mathcal{P}$) is clearly feasible and convex. The dual problem ($\mathcal{D}$) is also a convex quadratic program. Moreover, define $\alpha_{i}=\ell(\nu+\mu)/m$ if $i\in[1,m]$; Otherwise $\alpha_{i}=\ell\mu/n$. Then this ${\bm{\alpha}}$ is feasible to ($\mathcal{D}$) since $\mu\leq\min\{m/{\ell},bn/{\ell}\}$ and $\mu+\nu<m/\ell$. Owing to the well-known Frank-Wolfe theorem Frank and Wolfe (1956), both the quadratic programming problem ($\mathcal{P}$) as well as its dual problem ($\mathcal{D}$) are solvable. By (Bertsekas, 1999, Proposition 3.4.2 (a)), strong duality holds at the optimal primal-dual solution. The necessary and sufficient optimality conditions can be derived by applying the Karush-Kuhn-Tucker optimality conditions to ($\mathcal{P}$).
As for problem (2.7), its Lagrangian dual problem admits the following explicit form
$\begin{array}[]{ll}\mathop{\rm maximize}\limits_{{\bm{\alpha}}\in{\mathbb{R}}^{n}}&-\frac{1}{2l\lambda}\sum\limits_{i,j=m+1}^{\ell}k_{ij}\alpha_{i}\alpha_{j}-\frac{1}{2}\sum\limits_{i=m+1}^{\ell}\left(k_{ii}-\frac{2}{l\lambda}\sum\limits_{j=1}^{m}k_{ji}\right)\alpha_{i}-\frac{1}{2\ell\lambda}\sum\limits_{i,j=1}^{m}k_{ij}\\ {\mathrm{subject\ to}}&\sum\limits_{i=m+1}^{\ell}\alpha_{i}=\ell\mu,\\ &0\leq\alpha_{i}\leq b,i\in[m+1,\ell].\end{array}$ (4.2)
We show in the next theorem that the choice of parameters guarantee that that the optimal solutions of (2.7) always exist.
Theorem 8
Suppose $\mu+\nu\geq m/{\ell}$ and $\mu<m/{\ell}$ holds. Problem (2.7) and its dual problem (4.2) are solvable and the strong duality holds at the primal-dual optimal solution. In particular, $({\bm{a}},z,{\bm{\xi}},{\bm{\alpha}})$ is the primal-dual optimal solution if and only if it satisfies
${\bm{a}}=\frac{1}{\ell\lambda}\left[\sum_{i=1}^{m}\Phi({\bm{x}}_{i})+\sum_{i=m+1}^{\ell}y_{i}\alpha_{i}\Phi({\bm{x}}_{i})\right],$ (4.3a)
$\sum_{i=m+1}^{\ell}\alpha_{i}=\ell\mu,$ (4.3b)
$t-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}-2\xi_{i}\leq 0,\ \xi_{i}\geq 0,$ $\ i\in[m+1,\ell],$ (4.3c)
[…]
$(t-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}-2\xi_{i})\alpha_{i}=0,$ $\ i\in[m+1,\ell]$ (4.3e)
with $t=\|{\bm{a}}\|^{2}-z$.
Proof  First of all, $\lambda>0$ implies (2.7) is well defined and convex since $\mu<m/{\ell}$. Moreover, define $\alpha_{i}=\ell\mu/{n},i\in[m+1,\ell]$. Then this ${\bm{\alpha}}$ is feasible to (4.2) since (2.1). Owing to the well-known Frank-Wolfe theorem Frank and Wolfe (1956), both the quadratic programming problem (2.7) as well as its dual problem (4.2) are solvable. By (Bertsekas, 1999, Proposition 3.4.2 (a)), strong duality holds at the optimal primal-dual solution. The necessary and sufficient optimality conditions can be derived by applying the Karush-Kuhn-Tucker optimality conditions to (2.7).
As for problem (2.8), its Lagrangian dual problem is given by
$\begin{array}[]{cl}\mathop{{\mathrm{minimize}}}\limits_{\alpha\in{\mathbb{R}}^{l}}&\sum\limits_{i=m+1}^{\ell}k_{ii}\alpha_{i}\\ {\mathrm{subject\ to}}&\sum\limits_{i=m+1}^{\ell}\alpha_{i}=\ell\mu\\ &0\leq\alpha_{i}\leq b,\;i=m+1,\dots,\ell\\ &\sum\limits_{i=m+1}^{\ell}\alpha_{i}\Phi({\bm{x}}_{i})=\sum\limits_{i=1}^{m}\Phi({\bm{x}}_{i})\end{array}$ (4.4)
It is trivial that (2.8) is feasiable. If the dual problem (4.4) is infeasible, (2.8) is unbounded below, neither does problem ($\mathcal{F}$). A necessary condition for (4.4) to be feasible is $\mu/b\leq n/\ell$. However, we show this condition is not sufficient for the existence of the optimal solution to (2.8) with a counterexample.
Example 1
Let $m=n=1,\ell=2,\mu=1/2,b=1$, ${\bm{x}}_{1}=(1,0)$ with label $y_{1}=1$ and ${\bm{x}}_{2}=(0,1)$ with label $y_{2}=-1$. Considering the linear kernel, the dual problem (4.4) is infeasible in this case since $(0,\alpha_{2})\neq(1,0)$ for any $\alpha_{2}\in[0,1]$. Therefore, the corresponding (2.8) is unbounded below.
4.2 Uniqueness of the optimal sphere center
In this subsection, we characterize the optimal primal solution of the proposed ($\mathcal{P}$). In particular, we show that the global optimal solution is unique. Such a result represents the stark contrast of ($\mathcal{P}$), and is unrealistic to obtain for traditional methods such as (1.8) and (1.9) due to the lack of convexity.
First of all, the optimal center of the sphere is unique, since (2.5) and (2.7) is partially strictly convex with respect to ${\bm{a}}$.
Theorem 9
The optimal ${\bm{a}}_{*}$ for ($\mathcal{P}$) and (2.7) is unique if $m/\ell>\mu$.
Proof  It suffices to show that for any optimal solutions $({\bm{a}}^{\prime},s^{\prime},t^{\prime})$ and $({\bm{a}}^{\prime\prime},s^{\prime\prime},t^{\prime\prime})$ for (2.5), it holds that ${\bm{a}}^{\prime}={\bm{a}}^{\prime\prime}$. Suppose by contradiction that ${\bm{a}}^{\prime}\neq{\bm{a}}^{\prime\prime}$. By convexity,
${\bm{u}}:=\tfrac{1}{2}({\bm{a}}^{\prime},s^{\prime},t^{\prime})+\tfrac{1}{2}({\bm{a}}^{\prime\prime},s^{\prime\prime},t^{\prime\prime})$
is also optimal for (2.5). It follows from $\|\tfrac{1}{2}{\bm{a}}^{\prime}+\tfrac{1}{2}{\bm{a}}^{\prime\prime}\|^{2}<\tfrac{1}{2}\|{\bm{a}}^{\prime}\|^{2}+\tfrac{1}{2}\|{\bm{a}}^{\prime\prime}\|^{2}$ that
$f({\bm{u}})<\tfrac{1}{2}f({\bm{a}}^{\prime},s^{\prime},t^{\prime})+\tfrac{1}{2}f({\bm{a}}^{\prime\prime},s^{\prime\prime},t^{\prime\prime})=f({\bm{a}}^{\prime},s^{\prime},t^{\prime}),$
contradicting with the optimality of $({\bm{a}}^{\prime},s^{\prime},t^{\prime})$. Therefore, the optimal ${\bm{a}}_{*}$ for ($\mathcal{P}$) is unique. It follows from the same argument that ${\bm{a}}_{*}$ is also unique for (2.7) if $m/\ell>\mu$.
4.3 Uniqueness of the optimal radius and margin
Since the sphere center is unique for ($\mathcal{P}$), now we fix the sphere center and investigate the uniqueness of the radius and margin. For this purpose, notice that the strong duality holds for ($\mathcal{P}$) and ($\mathcal{D}$). A direct result from this property is that for any optimal primal variable $({\bm{a}},s,t,{\bm{\xi}})$ for ($\mathcal{P}$) and any optimal dual variable ${\bm{\alpha}}$ for ($\mathcal{D}$), $({\bm{a}},s,t,{\bm{\xi}},{\bm{\alpha}})$ together satisfies the KKT conditions (4.1). Therefore, our analysis is based on the optimality conditions (4.1) and the primal optimal solution is characterized via any given optimal dual variable.
Let ${\bm{\alpha}}^{*}$ be an optimal solution to the dual problem ($\mathcal{D}$). Define the following index sets
${\mathcal{T}}_{+}:=$$\{i:y_{i}=1,0\leq\alpha_{i}^{*}<1\},\quad{\mathcal{T}}_{-}:=\{i:y_{i}=-1,0\leq\alpha_{i}^{*}<b\},$
${\mathcal{SV}}_{+}:=$$\{i:y_{i}=1,0<\alpha_{i}^{*}\leq 1\},\ {\mathcal{SV}}_{-}:=\{i:y_{i}=-1,0<\alpha_{i}^{*}\leq b\}.$
Proposition 10
Suppose $\mu+\nu<m/\ell$ holds. Let ${\bm{\alpha}}^{*}$ be an optimal solution to the dual problem ($\mathcal{D}$). The unique optimal sphere center for ($\mathcal{P}$) is
${\bm{a}}_{*}=\frac{1}{\ell\nu}\sum\limits_{i=1}^{\ell}y_{i}\Phi({\bm{x}}_{i})\alpha_{i}^{*}.$ (4.5)
The optimal radii and margin for ($\mathcal{P}$) is
$\Gamma_{*}=\{(r,t):r_{l}\leq r\leq r_{u},q_{l}\leq r+t\leq q_{u},t\geq 0,t(\sum_{i=m+1}^{\ell}\alpha_{i}^{*}-\ell\mu)=0\}$ (4.6)
with $r_{l}:=\max\limits_{i\in{\mathcal{T}}_{+}}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}$, $r_{u}:=\min\limits_{i\in{\mathcal{SV}}_{+}}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}$, $q_{l}:=\max\limits_{i\in{\mathcal{SV}}_{-}}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}$ and $q_{u}:=\min\limits_{i\in{\mathcal{T}}_{-}}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}$, where by custom $\max\emptyset=-\infty$ and $\min\emptyset=+\infty$.
Proof  By duality, the primal-dual pair $({\bm{a}},s,t,{\bm{\xi}},{\bm{\alpha}}^{*})$ is optimal if and only if it satisfies the optimality conditions in (4.1) by Theorem 7. Hence, the vector ${\bm{a}}_{*}$ in (4.5) is optimal and unique due to (4.1c) and Theorem 9.
Since ${\bm{a}}_{*}$ is unique, we can treat it as fixed value and let $r=\|{\bm{a}}_{*}\|^{2}-s$ and $q=r+t$. The complementary slackness condition (4.1f) implies that $\xi_{i}=0$ for $i\in{\mathcal{T}}_{+}\cup{\mathcal{T}}_{-}$, meaning ${\bm{x}}_{i},i\in{\mathcal{T}}_{+}\cup{\mathcal{T}}_{-}$ are separated correctly by the sphere. The complementary slackness conditions (4.1g) and (4.1h) imply $\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}-r-2\xi_{i}=0,i\in{\mathcal{SV}}_{+}$ and $q-\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}-2\xi_{i}=0,i\in{\mathcal{SV}}_{-}$. Combining those conditions with the primal feasible conditions (4.1a) and (4.1b), $(r,t)$ are optimal if and only if $t\geq 0$ and the following conditions are satisfied:
$r\geq\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\text{ with }\xi_{i}=0,i\in{\mathcal{T}}_{+},$
$r\leq\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\text{ with }2\xi_{i}=\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}-r,i\in{\mathcal{SV}}_{+},$
$q\leq\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\text{ with }\xi_{i}=0,i\in{\mathcal{T}}_{-},$
$q\geq\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\text{ with }2\xi_{i}=q-\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2},i\in{\mathcal{SV}}_{-}.$
Then $(r,t)$ is optimal if and only if $(r,t)\in\Gamma_{*}$.
Remark 11
Note that $r$ always lies in a nonnegative finite interval. First, the dual feasibility $\sum_{i=1}^{\ell}y_{i}\alpha_{i}^{*}=\ell\nu$ and $\nu>0$ imply ${\mathcal{SV}}_{+}\neq\emptyset$. Hence $r_{u}\geq 0$. Secondly, if ${\mathcal{T}}_{+}\neq\emptyset$ then $r_{l}\geq 0$. Otherwise, $\alpha_{i}^{*}=1$ for all $y_{i}=1$. By the equality in (4.1d) one has $\sum_{i=m+1}^{\ell}\alpha_{i}^{*}=m-\ell\nu$. Combining it with $\mu+\nu<m/\ell$, $\sum_{i=m+1}^{\ell}\alpha_{i}^{*}>\ell\mu$, we have ${\mathcal{SV}}_{-}\neq\emptyset$ and $t_{*}=0$ by the complementarity in (4.1i). Hence $q_{l}\geq 0$ and $r\geq\max\{r_{l},q_{l}\}$.
We can then use the above characterization of $(r,t)$ to derive the following results on the uniqueness of $(r,t)$.
Theorem 12
Suppose $\mu+\nu<m/\ell$ holds. If there exist $i_{+}\in[1,m]$ with $\alpha^{*}_{i_{+}}\in(0,1)$ and $i_{-}\in[m+1,\ell]$ with $\alpha^{*}_{i_{-}}\in(0,b)$, then $r_{*}=\|\Phi({\bm{x}}_{i_{+}})-{\bm{a}}_{*}\|^{2}\ \text{ and }\ t_{*}=\|\Phi({\bm{x}}_{i_{-}})-{\bm{a}}_{*}\|^{2}-\|\Phi({\bm{x}}_{i_{+}})-{\bm{a}}_{*}\|^{2}$ are the unique optimal radius and margin for ($\mathcal{P}$), respectively. If $\sum_{i=m+1}^{\ell}\alpha_{i}^{*}>\ell\mu$ then the optimal margin is 00 and the optimal radii set is
$\left[\max\{r_{l},q_{l}\},\min\{r_{u},q_{u}\}\right].$ (4.7)
In particular, if there exist $i_{*}\in[1,m]$ with $\alpha^{*}_{i_{*}}\in(0,1)$ or $i_{*}\in[m+1,\ell]$ with $\alpha^{*}_{i_{*}}\in(0,b)$, then $r_{*}=\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}$ is the unique optimal radius for ($\mathcal{P}$).
Proof  Under the assumption, one has $i_{+}\in{\mathcal{T}}_{+}\cap{\mathcal{SV}}_{+}$ and $i_{-}\in{\mathcal{T}}_{-}\cap{\mathcal{SV}}_{-}$. Then the optimal radii and margin for ($\mathcal{P}$) in (4.6) reduces to the single point
$\Gamma_{*}=\{(\|\Phi({\bm{x}}_{i_{+}})-{\bm{a}}_{*}\|^{2},\|\Phi({\bm{x}}_{i_{-}})-{\bm{a}}_{*}\|^{2}-\|\Phi({\bm{x}}_{i_{+}})-{\bm{a}}_{*}\|^{2})\}.$
If $\sum_{i=m+1}^{\ell}\alpha_{i}^{*}>\ell\mu$ then the optimal margin is 00 by the complementary condition (4.1i). Bring $t=0$ into (4.6), one has the desired optimal radii set (4.7).
If there exist $i_{*}\in[1,m]$ with $\alpha^{*}_{i_{*}}\in(0,1)$ then $i_{*}\in{\mathcal{T}}_{+}\cap{\mathcal{SV}}_{+}$. Hence the interval in (4.7) reduces to the point $\|\phi({\bm{x}}_{i_{*}})-{\bm{a}}_{*}\|^{2}$. Similarly if there exists $i_{*}\in[m+1,\ell]$ with $\alpha^{*}_{i_{*}}\in(0,b)$, then $i_{*}\in{\mathcal{T}}_{-}\cap{\mathcal{SV}}_{-}$. Hence the interval in (4.7) also reduces to the point $\|\phi({\bm{x}}_{i_{*}})-{\bm{a}}_{*}\|^{2}$. This completes the proof.
As for problem (2.7), we can also have the following similar result.
Proposition 13
Suppose $0<\mu<m/{\ell}$ and $\mu+\nu\geq m/\ell$ hold. Let ${\bm{\alpha}}^{*}$ be an optimal solution to the dual problem (4.2). The unique optimal sphere center for (2.7) is
${\bm{a}}_{*}=\frac{1}{\ell\lambda}\left(\sum_{i=1}^{m}\Phi({\bm{x}}_{i})+\sum_{i=m+1}^{\ell}y_{i}\Phi({\bm{x}}_{i})\alpha_{i}^{*}\right),$ (4.8)
and the optimal radius is $r_{*}=0$. If $\sum_{i=1}^{\ell}\alpha_{i}^{*}>\ell\mu$ then the optimal margin is 00. Otherwise, the optimal margin set is
$\Gamma_{*}=\{t:\max\limits_{i\in{\mathcal{SV}}_{-}}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\leq t\leq\min\limits_{i\in{\mathcal{T}}_{-}}\|\Phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\}.$ (4.9)
If there exist $i_{-}\in[m+1,\ell]$ with $\alpha^{*}_{i_{-}}\in(0,b)$, then $t_{*}=\|\Phi({\bm{x}}_{i_{-}})-{\bm{a}}_{*}\|^{2}$ is the unique optimal margin for (2.7).
Proof  First of all, Theorem 8 and Theorem 4 implies (2.7) is solvable and the optimal radius is $r_{*}=0$. By duality, the primal-dual pair $({\bm{a}}_{*},t,{\bm{\xi}},{\bm{\alpha}}^{*})$ is optimal if and only if it satisfies the optimality conditions in (4.3). Hence, the vector ${\bm{a}}_{*}$ in (4.8) is optimal and unique due to Theorem 9.
Since ${\bm{a}}_{*}$ is given and $r_{*}=0$, it follows from the optimality condition in (4.3) that $\xi_{i}=0$ and $\|\phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\leq t$ for any $i\in{\mathcal{T}}_{-}$ and $\|\phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}-t-2\xi_{i}=0$ for any $i\in{\mathcal{SV}}_{-}$. Therefore, $t$ is optimal if and only if $t\geq 0$ and the following conditions are satisfied:
$t\leq\|\phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\text{ with }\xi_{i}=0,\qquad i\in{\mathcal{T}}_{-},$ (4.10)
$t\geq\|\phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2}\text{ with }2\xi_{i}=t-\|\phi({\bm{x}}_{i})-{\bm{a}}_{*}\|^{2},\ i\in{\mathcal{SV}}_{-}.$ (4.11)
Then $t$ is optimal if and only if $(r,t)\in\Gamma_{*}$. Under the assumption, one has $i_{-}\in{\mathcal{T}}_{-}\cap{\mathcal{SV}}_{-}$. Then the optimal margin set $\Gamma_{*}$ in (4.9) reduces to the single point $\{\|\phi({\bm{x}}_{i_{-}})-{\bm{a}}_{*}\|^{2}\}$.
4.4 $\nu$-property
Following the terminology in Schölkopf et al. (2001); Scholkopf et al. (2000); Wu and Ye (2009), a training example ${\bm{x}}_{i},1\leq i\leq\ell$ is called a support vector (SV) if the associated multiplier $\alpha^{*}_{i}>0$. Therefore, $\{{\bm{x}}_{i}\mid i\in{\mathcal{SV}}_{-}\cup{\mathcal{SV}}_{+}\}$ are the set of SVs. ${\bm{x}}_{i}$ is a margin error (ME) if the corresponding $\xi^{*}_{i}>0$. As for MEs, by complementary condition (4.1f),
$\alpha_{i}^{*}=\begin{cases}1&\text{ if ${\bm{x}}_{i}$ is a positive ME},\\ b&\text{ if ${\bm{x}}_{i}$ is a negative ME}.\end{cases}$ (4.12)
Therefore, the MEs in the positive constitute a subset of ${\mathcal{SV}}_{+}$ and the MEs in the negative is a subset of ${\mathcal{SV}}_{-}$. Let $m_{+}$ and $n_{+}$ denote the number of MEs in the positive and negative classes, while $s_{+}$ and $s_{-}$ denote the number of SVs in the positive and negative classes, respectively. This means
$m_{+}\leq|{\mathcal{SV}}_{+}|,\quad n_{+}\leq|{\mathcal{SV}}_{-}|,\quad s_{+}=|{\mathcal{SV}}_{+}|,\quad s_{-}=|{\mathcal{SV}}_{-}|.$
The $\nu$-property analyzed in OCSVM Schölkopf et al. (1999, 2001), $\nu$-SVM Scholkopf et al. (2000) and the traditional SSLM Wu and Ye (2009) established the connection between the model parameters and the fractions of SV and ME in the dataset. Our model ($\mathcal{P}$) involves a nonnegative parameter $\mu$ and two positive parameters $\nu$ and $b$. We can also derive a similar $\nu$-property for these parameters in the following proposition.
Proposition 14
Suppose $\mu+\nu<m/\ell$ holds. If the optimal margin of ($\mathcal{P}$) is positive, then
$\frac{m_{+}}{\ell}\leq\nu+\mu\leq\frac{s_{+}}{\ell},\quad\frac{n_{+}}{\ell}\leq\frac{\mu}{b}\leq\frac{s_{-}}{\ell};$ (4.13)
Otherwise,
$\frac{m_{+}}{\ell}\leq\nu+b\frac{s_{-}}{\ell},\quad\nu+b\frac{n_{+}}{\ell}\leq\frac{s_{+}}{\ell},\quad\frac{\mu}{b}\leq\frac{s_{-}}{\ell}.$ (4.14)
Proof  From the optimality condition (4.1d), there exists an optimal solution ${\bm{\alpha}}^{*}$ of ($\mathcal{D}$) satisfies
$\sum_{i=1}^{m}\alpha^{*}_{i}-\sum_{i=m+1}^{\ell}\alpha^{*}_{i}=\ell\nu.$ (4.15)
and
$\sum_{i=m+1}^{\ell}\alpha^{*}_{i}\geq\ell\mu.$ (4.16)
If $t_{*}>0$, by (4.1i), it holds that
$\sum_{i=m+1}^{\ell}\alpha_{i}^{*}=\ell\mu$ (4.17)
and
$\sum_{i=1}^{m}\alpha_{i}^{*}=\ell(\nu+\mu).$ (4.18)
Summing up $\alpha^{*}_{i}$ over the positive MEs leads to
$m_{+}=\sum_{{\bm{x}}_{i}\text{ is positive ME}}\alpha_{i}^{*}\leq\sum_{i\in{\mathcal{SV}}_{+}}\alpha_{i}^{*}=\sum_{i=1}^{m}\alpha_{i}^{*},$ (4.19)
where the last equality is by the fact that $\alpha_{i}^{*}=0,i\in[1,m]\setminus{\mathcal{SV}}_{+}$. One the other hand,
$s_{+}=|{\mathcal{SV}}_{+}|\geq\sum_{i\in{\mathcal{SV}}_{+}}\alpha_{i}^{*}=\sum_{i=1}^{m}\alpha_{i}^{*}.$ (4.20)
Combining (4.18), (4.19) and (4.20) we can derive the first group inequalities in (4.13).
Summing up $\alpha^{*}_{i}$ over the negative MEs leads to
$bn_{+}=\sum_{{\bm{x}}_{i}\text{ is negative ME}}\alpha_{i}^{*}\leq\sum_{i\in{\mathcal{SV}}_{-}}\alpha_{i}^{*}=\sum_{i=m+1}^{\ell}\alpha_{i}^{*},$ (4.21)
where the last equality is by the fact that $\alpha_{i}^{*}=0,i\in[m+1,\ell]\setminus{\mathcal{SV}}_{-}$. On the other hand,
$bs_{-}\geq\sum_{i\in{\mathcal{SV}}_{-}}\alpha_{i}^{*}=\sum_{i=m+1}^{\ell}\alpha_{i}^{*}.$ (4.22)
The second group inequalities in (4.13) is then derived based on (4.17), (4.21) and (4.22).
If the margin $t_{*}=0$, it follows from (4.19), (4.22) and (4.15) that
$m_{+}-bs_{-}\leq\sum_{i=1}^{m}\alpha^{*}_{i}-\sum_{i=m+1}^{\ell}\alpha^{*}_{i}=\ell\nu.$ (4.23)
Similarly, it follows from (4.20), (4.21) and (4.15) that
$s_{+}-bn_{+}\geq\sum_{i=1}^{m}\alpha^{*}_{i}-\sum_{i=m+1}^{\ell}\alpha^{*}_{i}=\ell\nu.$ (4.24)
Combining (4.23) and (4.24) yields the first two inequalities in (4.14). The last inequality in (4.14) is true since by (4.16) one has
$bs_{-}\geq\sum_{i\in{\mathcal{SV}}_{-}}\alpha^{*}_{i}=\sum_{i=m+1}^{\ell}\alpha^{*}_{i}\geq\ell\mu,$
completing the proof.
In the absence of the negative data where $\mu=0$ and all the terms regarding the negative data disappear in ($\mathcal{P}$), then assumption $\mu+\nu<m/\ell$ reduces to $\nu<1$ and one can easily arrive at the following result for this particular case.
Corollary 15
Assume $\nu\in(0,1)$ and $m=\ell$. Then the solution of ($\mathcal{P}$) satisfies
$\frac{m_{+}}{m}\leq\nu\leq\frac{s_{+}}{m},$ (4.25)
that is $\nu$ is an upper bound on the fraction of errors, and is a lower bound on the fraction of SVs.
The following result for the degenerate case can also be easily derived by a similar argument based on (4.3b).
Proposition 16
Suppose $\mu<m/{\ell}$ and $\mu+\nu\geq m/\ell$ holds. One has
$\frac{n_{+}}{\ell}\leq\frac{\mu}{b}\leq\frac{s_{-}}{\ell}.$ (4.26)
Proof  Summing up $\alpha^{*}_{i}$ over the negative MEs leads to
$bn_{+}=\sum_{{\bm{x}}_{i}\text{ is negative ME}}\alpha_{i}^{*}\leq\sum_{i\in{\mathcal{SV}}_{-}}\alpha_{i}^{*}\leq\sum_{i=m+1}^{\ell}\alpha_{i}^{*}=\ell\mu$ (4.27)
by (4.3b) and (4.12). One the other hand,
$bs_{-}\geq\sum_{i\in{\mathcal{SV}}_{-}}\alpha_{i}^{*}=\sum_{i=m+1}^{\ell}\alpha_{i}^{*}=\ell\mu.$ (4.28)
Combining (4.27) and (4.28) we can derive (4.26).
5 Numerical Experiments
In this section, we present the experimental results to demonstrate the effectiveness of our proposed method. The comparison experiments were conducted on the Banana dataset [^2] https://leon.bottou.org/papers/bordes-ertekin-weston-bottou-2005, a synthetic two-dimensional dataset containing 5,300 data points organized into two banana-shaped clusters. Approximately 10% of the dataset represents a noisy region where positive and negative data points cannot be distinctly separated. To reduce noise in the dataset, we assigned a label of $-1$ to data points satisfying the condition $x_{2}+\frac{3}{7}(x_{1}-3)>0$.
Table 1: Summary of the Banana data sets for five Scenarios and test set.
Scenario A Scenario B Scenario C Scenario D Scenario E Test
$(m,n)$ $(3,3)$ $(100,100)$ (768,1232) (1000,100) (1000,0) (381,619)
We consider five scenarios in our analysis. Table 1 summarizes the training datasets across these five scenarios, along with a test dataset. Scenarios A, B, and C are designed for binary classification, whereas Scenarios D and E focus on anomaly detection, using supervised and clean training data, respectively. Notably, Scenario D shares positive data points with Scenario E.
We investigate the effect of the nonconvexity in state--of-the--art models and compare the classification boundaries learned by our proposed primal CSSLM ($\mathcal{P}$), dual CSSLM ($\mathcal{D}$), primal SSLM (1.9), dual SSLM (1.10), SVDD, SVM (1.7) and OCSVM (1.6). The experiments were executed on a machine with an Intel Core i7-13700H CPU and 16GB of RAM.
5.1 Comparison of CSSLM with SVM and OCSVM
We first examine the boundaries with Gaussian Kernel learned by dual CSSLM (setting $\mu=0$) on Scenario C, D and E, SVM on Scenario C and D, and OCSVM on Scenario E, a one-class dataset. Parameters $\nu$ and $\gamma$ (controlling the Gaussian kernel width $k({\bm{x}},{\bm{y}})=e^{-\gamma\|{\bm{x}}-{\bm{y}}\|^{2}}$) were selected through cross-validation. For CSSLM and SVM, we set $b=1$ in Scenario C and $b=10$ in Scenario D. Dual CSSLM was solved using quadprog in Matlab, while OCSVM and SVM were solved using libsvm in Matlab. Table 2 displays the parameters used and the classification errors on the test data, while Figure 4 illustrates the boundaries produced by each model. Figure 5 shows an overlay of the boundaries learned by CSSLM and its competitors for direct comparison.
Table 2: $\nu$ and $\gamma$ used in three models and the comparison of the dual CSSLM on all scenarios with SVM on Scenario C and D, and OCSVM on Scenario E. Test error on positive class is $e_{+}$ and test error on negative class is $e_{-}$.
CSSLM SVM/OCSVM
$\nu$ $\gamma$ $e_{+}(\%)$ $e_{-}(\%)$ $\nu$ $\gamma$ $e_{+}(\%)$ $e_{-}(\%)$
Scenario C $10^{-2}$ $0.2$ $11.02$ $5.49$ $10^{-4}$ $2.0$ $9.19$ $4.20$
Scenario D $0.5$ $1.0$ $54.33$ $0.16$ $10^{-5}$ $0.1$ $10.24$ $4.52$
Scenario E $0.5$ $1.0$ $49.87$ $0.48$ $0.5$ $1.0$ $49.87$ $0.48$
(a) Scenario C
(b) Scenario D
(c) Scenario E
Figure 4: Boundaries learned by CSSLM, SVM and OCSVM in three scenarios and the training data in which the data in circle is the supporting vector.
(a) CSSLM
(b) SVM/OCSVM
Figure 5: Test data and boundaries learned by CSSLM, SVM, OCSVM in three scenarios.
Observations:
  • In Scenarios C and D, CSSLM consistently identifies the banana-shaped boundary, whereas SVM with a Gaussian kernel overfits, producing an aggressive boundary around positive data, which explains its lower positive errors. In Scenario E, CSSLM and OCSVM yield nearly identical banana-shaped boundaries, underscoring CSSLM’s adaptability to various data types.
  • Figure 5 highlights CSSLM’s resilience and strong generalization across datasets, as it consistently identifies similar banana-shaped boundaries.
5.2 Nonconvexity issues in solving the primal problem
To illustrate challenges in solving the primal SSLM due to nonconvexity, we applied several advanced nonlinear solvers available in AMPL for the 10-dimensional feature space in Scenarios A, B, and C using the transformation:
$\Phi({\bm{x}})=(1,\sqrt{3}x_{1},\sqrt{3}x_{2},\sqrt{3}x_{1}^{2},\sqrt{3}x_{2}^{2},\sqrt{6}x_{1}x_{2},x_{1}^{3},x_{2}^{3},\sqrt{3}x_{1}x_{2}^{2},\sqrt{3}x_{1}^{2}x_{2}),$
which corresponds to the cubic polynomial kernel function $k({\bm{x}},{\bm{y}})=(1+{\bm{x}}^{T}{\bm{y}})^{3}$. These solvers include the global optimization solver Gurobi and three local optimization solvers—Ipopt, Knitro, and Snopt.
For consistency in comparison, the optimal values of SSLM and dual SSLM were scaled by a factor of $\ell\nu/2$ to match the optimal values of the primal CSSLM problem ($\mathcal{P}$) and the dual CSSLM problem ($\mathcal{D}$). Parameter choices were $b=1,\nu=0.001$ and $\mu=0.05$, ensuring $\mu+\nu<m/\ell$. The SSLM parameters $(\bar{\nu},\bar{\nu}_{1},\bar{\nu}_{2})$ were configured based on Equation (3.2). Table 3 presents the optimal values, exit flags [^3] Exitflag$=$1 means the solver exits normally. Exitflag$=$2 means Restoration Failed for Ipopt, which indicates that Ipopt failed to find a feasible point in the feasibility restoration phase., runtime, and optimality residuals obtained by each solver.
Table 3: The optimal values and exit flag obtained, runtimes and optimality errors of Gurobi, Ipopt, Knitro and Snopt for the primal (1.9) in 10-dimensional feature space of Scenario A, B and C .
Scenario Solver Optimal value Exitflag Runtime Optimality Error
A Gurobi ${\bf-225.8368}$ $1$ $4.7900$ 9.4735e--9
Ipopt $1.3246$ $1$ $11.47$ 4.5793e--7
Knitro $1.3246$ $1$ $0.04$ 1.1400e--15
Snopt $1.3246$ $1$ $0.00$ 7.0000e--7
B Gurobi ${\bf 121.9522}$ $1$ $0.96$ 1.1057e--8
Ipopt $518.9560$ $2$ $14.09$ 1.5224e$+$3
Knitro $1586.8000$ $1$ $114.24$ 3.3000e--12
Snopt $1586.8000$ $1$ $0.03$ 1.3000e--7
C Gurobi ${\bf 1747.4000}$ $1$ $13.13$ 6.0332e--9
Ipopt $5981.3600$ $2$ $1532.11$ 3.3145e$+$4
Knitro $12024.7000$ $1$ $1666.35$ 1.3200e--11
Snopt $12024.7000$ $1$ $2.76$ 5.8000e--7
Observations:
  • In all three scenarios, Gurobi successfully identified the global optimal solution (highlighted in bold), but its computational time was higher than that of successful local solvers such as Snopt.
  • Notably, none of the local solvers consistently identified the global optimum. Ipopt, for example, failed to find a feasible solution in Scenarios B and C. Additionally, some local solvers, such as Knitro and Ipopt, required significantly more computation time (up to 100 times longer than Gurobi). A large discrepancy between the local and global optimal values indicates that local solutions may diverge considerably from the true global optimum.
5.3 Nonconvexity issues in solving the dual problem
As demonstrated, achieving the global optimum for the nonconvex primal SSLM is challenging. A natural alternative is to solve the convex dual SSLM (1.10) instead of the primal problem. However, our experiment shows that the dual SSLM does not always yield the global optimal solution, reflecting the lack of strong duality.
We used quadprog in Matlab to solve the dual SSLM (1.10), primal CSSLM ($\mathcal{P}$), and its dual problem ($\mathcal{D}$) in Scenarios A, B, and C, all of which are convex QPs. Table 4 lists the optimal objectives, exit flags444exitflag$=$1 in quadprog means the optimal solution is found, runtime, and optimality errors.
Table 4: The optimal values, obtained, runtimes and optimality residuals of solvers for the dual SSLM (1.10), proposed problem ($\mathcal{P}$) and its dual problem ($\mathcal{D}$) in 10-dimensional feature space of Scenario A, B and C.
Scenario Model Optimal value Exitflag Runtime Optimality Error
A Dual SSLM (1.10) $-225.8368$ $1$ $0.68$ 4.5475e--13
CSSLM ($\mathcal{P}$) $-225.8368$ $1$ $0.10$ 4.1457e--9
CSSLM ($\mathcal{D}$) $-225.8368$ $1$ $0.02$ 3.4494e--12
B Dual SSLM (1.10) ${\bf 121.7294}$ $1$ $0.75$ 1.7144e--11
CSSLM ($\mathcal{P}$) $121.9522$ $1$ $0.09$ 6.5391e--11
CSSLM ($\mathcal{D}$) $121.9522$ $1$ $0.05$ 3.1775e--10
C Dual SSLM (1.10) ${\bf 1733.1000}$ $1$ $6.41$ 4.1751e--9
CSSLM ($\mathcal{P}$) $1747.4000$ $1$ $16.02$ 1.6268e--10
CSSLM ($\mathcal{D}$) $1747.4000$ $1$ $12.43$ 4.2415e--9
Observations:
  • The dual optimal solution for SSLM does not always correspond to the global optimal solution of the primal SSLM, as observed in Scenarios B and C (in bold). This highlights the absence of strong duality in nonconvex SSLM, indicating that solving the dual SSLM does not resolve nonconvexity issues present in the primal problem.
  • Our proposed primal and dual CSSLM consistently achieved the global optimal solution, which confirms strong duality in our model, a clear advantage over nonconvex alternatives. Additionally, in Scenarios A and B, quadprog was substantially faster when solving CSSLM for the global solution.
6 Conclusion
We proposed a novel convex SSLM formulation, which simplifies to a convex quadratic programming problem for specific hyperparameter values. The convexity of our method allows us to derive several valuable results that are not achievable with traditional nonconvex SSLM models, greatly simplifying the analysis of hypersphere--based SVM solutions. We thoroughly examined how hyperparameters influence the optimal solution and established clear connections between our model and traditional SSLM, specifying conditions under which the optimal solution is unique. We also derive the $\nu$-property, which explains how hyperparameters affect the distribution of support vectors and margin errors across positive and negative classes. The numerical experiments confirmed that our proposed model outperforms traditional primal SSLM and its dual in terms of both efficiency and accuracy.
We believe that resolving the issues associated with nonconvexity will drive further advancements in developing robust algorithms for hypersphere--based SVMs. Future work will explore additional properties and algorithms stemming from the convexity of our model.
Acknowledgment This work was initiated from a discussion between the first author and Shuisheng Zhou at Xidian University on the nonconvexity in anomaly detection. The authors gratefully acknowledge the many stimulating suggestions and comments from Shuisheng Zhou during the preparation of this paper.
Hongying Liu is supported by the National Natural Science Foundation of China under Grant 12171021 and 12131004.
Appendix A Proof of Theorem 1 in the ill-posed cases
Proof  If $\mu>m/{\ell}$, setting $({\bm{a}},r,t)=(\alpha\tilde{\bm{a}},0,\alpha^{2}\|\tilde{\bm{a}}\|^{2})$ with any $\alpha>0$ and $\tilde{\bm{a}}\neq\bm{0}$ yields
$\begin{array}[]{rl}&g(\alpha\tilde{\bm{a}},0,\alpha^{2}\|\tilde{\bm{a}}\|^{2})\\ =&-\frac{1}{2}\mu\alpha^{2}\|\tilde{\bm{a}}\|^{2}+\frac{1}{2\ell}\sum\limits_{i=1}^{m}\|\Phi({\bm{x}}_{i})-\alpha\tilde{\bm{a}}\|^{2}+\frac{b}{2\ell}\sum\limits_{i=m+1}^{\ell}\left(\alpha^{2}\|\tilde{\bm{a}}\|^{2}-\|\Phi({\bm{x}}_{i})-\alpha\tilde{\bm{a}}\|^{2}\right)_{+}\\ =&\frac{1}{2}(\tfrac{m}{\ell}-\mu)\alpha^{2}\|\tilde{\bm{a}}\|^{2}-\tfrac{1}{\ell}\sum\limits_{i=1}^{m}\langle\Phi({\bm{x}}_{i}),\tilde{\bm{a}}\rangle\alpha+\frac{b}{\ell}\sum\limits_{i=m+1}^{\ell}\left(\langle\Phi({\bm{x}}_{i}),\tilde{\bm{a}}\rangle\alpha-\frac{k_{ii}}{2}\right)_{+}+\tfrac{1}{2\ell}\sum\limits_{i=1}^{m}k_{ii}\\ \leq&\frac{1}{2}(\tfrac{m}{\ell}-\mu)\alpha^{2}\|\tilde{\bm{a}}\|^{2}-\tfrac{1}{\ell}\left[\sum\limits_{i=1}^{m}\langle\Phi({\bm{x}}_{i}),\tilde{\bm{a}}\rangle-b\sum\limits_{i=m+1}^{\ell}\left(\langle\Phi({\bm{x}}_{i}),\tilde{\bm{a}}\rangle\right)_{+}\right]\alpha+\tfrac{1}{2\ell}\sum\limits_{i=1}^{m}k_{ii},\end{array}$
where the inequality is by $k_{ii}\geq 0$. It follows that $g(\alpha\tilde{\bm{a}},0,\alpha^{2}\|\tilde{\bm{a}}\|^{2})\to-\infty$ as $\alpha\to+\infty$, since $\mu>m/{\ell}$.
If $\mu>bn/{\ell}$, setting $({\bm{a}},r,t)=(0,0,t)$ with sufficiently large $t>0$ yields
$g(0,0,t)=\frac{1}{2}\Big{(}\frac{bn}{\ell}-\mu\Big{)}t+\frac{1}{2\ell}\sum_{i=1}^{m}\|\Phi({\bm{x}}_{i})\|^{2}-\frac{b}{2\ell}\sum_{i=m+1}^{\ell}\|\Phi({\bm{x}}_{i})\|^{2},$
which implies $g(0,0,t)\to-\infty$ as $t\to+\infty$.
Appendix A Proof of Theorem 3
Proof  (i) Substituting $r=0$ into $g(a,r,t)$, one has
$\begin{array}[]{ll}g({\bm{a}},r,t)&=\frac{\nu}{2}r+\frac{1}{2\ell}\sum\limits_{i=1}^{m}(\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}-r)_{+}+\frac{b}{2\ell}\sum\limits_{i=m+1}^{\ell}(r+t-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2})_{+}\\ &\geq\frac{1}{2}(\nu-\tfrac{m}{\ell})r+\frac{1}{2\ell}\sum\limits_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}\\ &\geq\frac{1}{2\ell}\sum\limits_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}\\ &\geq\frac{1}{2\ell}\sum\limits_{i=1}^{m}\|\Phi({\bm{x}}_{i})-\hat{\bm{a}}\|^{2}.\end{array}$
Here the first inequality holds as equality if and only if $r\leq\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2},i\in[1,m]$ and $r+t\leq\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2},i\in[m+1,\ell]$; the second inequality holds because of $\mu=0,\nu=m/\ell$ and the inequality holds as equality if and only if ${\bm{a}}=\hat{\bm{a}}$ and $(\nu-m/{\ell})r/2=0$. The last inequality is by the definition (2.3) of $\hat{\bm{a}}$. Therefore, $g({\bm{a}},r,t)$ reaches its lower bound if and only $({\bm{a}},r,t)\in\Omega$.
(ii) Suppose $\nu+\mu\geq m/{\ell}$ is satisfied. We first prove ($\mathcal{F}$) is equivalent to
$\inf\limits_{{\bm{a}}\in F,t\in\mathbb{R}}g({\bm{a}},0,t).$ (A.1)
Moreover, if $\inf\limits_{{\bm{a}}\in F,t\in\mathbb{R}}g({\bm{a}},0,t)>-\infty$, then the global minimizer satisfies $r=0$ and $t\geq 0$. By $\nu+\mu\geq m/{\ell}$, one has $0<\mu\leq m/{\ell}$. Hence for any ${\bm{a}}\in F,r\geq 0$,
$\begin{array}[]{rl}g({\bm{a}},r,t)\geq&\frac{1}{2}(\nu r-\mu t)+\frac{1}{2\ell}\sum_{i=1}^{m}(\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}-r)+\frac{b}{2\ell}\sum_{i=m+1}^{\ell}(r+t-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2})_{+}\\ =&\frac{1}{2}(\nu-\frac{m}{\ell})r-\frac{\mu}{2}t+\frac{1}{2\ell}\sum_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}+\frac{b}{2\ell}\sum_{i=m+1}^{\ell}(r+t-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2})_{+}\\ \geq&-\frac{\mu}{2}(r+t)+\frac{1}{2\ell}\sum_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}+\frac{b}{2\ell}\sum_{i=m+1}^{\ell}(r+t-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2})_{+}\\ =&g({\bm{a}},0,r+t)\\ \geq&\inf\limits_{{\bm{a}}\in F,t}g({\bm{a}},0,t)\end{array}$
where the first inequality comes from $(z)_{+}\geq z$ and the second inequality is due to condition $\nu+\mu\geq m/{\ell}$. On the other hand, for any $t<0$,
$g({\bm{a}},0,t)=-\frac{\mu}{2}t+\frac{1}{2\ell}\sum\limits_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}>\frac{1}{2\ell}\sum\limits_{i=1}^{m}\|\Phi({\bm{x}}_{i})-{\bm{a}}\|^{2}=g({\bm{a}},0,0)$ (A.2)
by $\mu>0$. It follows that $g({\bm{a}},r,t)\geq\inf\limits_{{\bm{a}}\in F,t}g({\bm{a}},0,t)\geq\inf\limits_{{\bm{a}}\in F,t\geq 0}g({\bm{a}},0,t)$. It implies
$\inf\limits_{{\bm{a}}\in F,t}g({\bm{a}},0,t)=\inf\limits_{{\bm{a}}\in F,t\geq 0}g({\bm{a}},0,t).$
Moreover, (A.2) indicates that if $\inf\limits_{{\bm{a}}\in F,t}g({\bm{a}},0,t)>-\infty$, $({\bm{a}}^{*},t^{*})\in\arg\min\limits_{{\bm{a}}\in F,t\in\mathbb{R}}g({\bm{a}},0,t)$ also satisfies $t^{*}\geq 0$. The former result implies that problem ($\mathcal{F}$) in this case is equivalent to the unconstrained problem (A.1), which is equivalent to a convex problem by setting $z=\|{\bm{a}}\|^{2}-t$ in $g$:
$\min\limits_{{\bm{a}}\in F,z\in{\mathbb{R}}}\tilde{f}({\bm{a}},z),$ (A.3)
with (see Appendix A for deriving $\tilde{f}({\bm{a}},z)$)
$\tilde{f}({\bm{a}},z)=\frac{\ell\lambda}{2}\|{\bm{a}}\|^{2}+\frac{\ell\mu}{2}z-m\langle\hat{\bm{a}},{\bm{a}}\rangle+b\sum_{i=m+1}^{\ell}\left(-\frac{z+k_{ii}}{2}+\langle{\bm{a}},\Phi({\bm{x}}_{i})\rangle\right)_{+}.$
If $\mu<m/\ell\leq\nu+\mu$, then (A.3) reverts to the convex quadratic program problem (2.7).
(iii) If $\mu=m/\ell$, then (A.3) reverts to the linear programming problem (2.8).
Appendix A Derivation of the objective function in the degenerate case
It holds that
$g({\bm{a}},0,t)=$ $-\frac{\mu}{2}t+\frac{1}{2\ell}\sum\limits_{i=1}^{m}\|{\bm{a}}-\Phi({\bm{x}}_{i})\|^{2}+\frac{b}{2\ell}\sum_{i=m+1}^{\ell}(t-\|\Phi({\bm{x}}_{i})-{\bm{a}}\|_{2}^{2})_{+}$
$=$ $-\frac{\mu}{2}t+\frac{1}{2\ell}\sum\limits_{i=1}^{m}(\|{\bm{a}}\|^{2}-2\langle{\bm{a}},\Phi({\bm{x}}_{i})\rangle+k_{ii})+\frac{b}{2\ell}\sum_{i=m+1}^{\ell}(t-\|{\bm{a}}\|^{2}+2\langle{\bm{a}},\Phi({\bm{x}}_{i})\rangle-k_{ii})_{+}$
$=$ $-\frac{\mu}{2}t+\frac{m}{2\ell}\|{\bm{a}}\|^{2}-\frac{m}{\ell}\langle\hat{\bm{a}},{\bm{a}}\rangle+\frac{b}{2\ell}\sum_{i=m+1}^{\ell}(t-\|{\bm{a}}\|^{2}+2\langle{\bm{a}},\Phi({\bm{x}}_{i})-k_{ii}\rangle)_{+}+\frac{1}{2\ell}\sum_{i=1}^{m}k_{ii}$
where $\hat{\bm{a}}=\frac{1}{m}\sum_{i=1}^{m}\Phi({\bm{x}}_{i})$. Letting $z=\|{\bm{a}}\|^{2}-t$ in $f$, it follows that
$g({\bm{a}},0,t)=\frac{1}{\ell}\tilde{f}({\bm{a}},z)+\frac{1}{2\ell}\sum_{i=1}^{m}k_{ii}$
where
$\tilde{f}({\bm{a}},z)=$ $\frac{\ell}{2}\left(\frac{m}{\ell}-\mu\right)\|{\bm{a}}\|^{2}+\frac{\ell\mu}{2}z-m\langle\hat{\bm{a}},{\bm{a}}\rangle+b\sum_{i=m+1}^{\ell}\left(-\frac{z+k_{ii}}{2}+\langle{\bm{a}},\Phi({\bm{x}}_{i})\rangle\right)_{+}.$